OpenStudy (anonymous):

Margaret traces a circle with a radius of 20 cm like the one shown below. She will color the shaded region. What is the approximate area of the shaded region?

OpenStudy (anonymous):



OpenStudy (anonymous):

smr_kml, could you please help me?

OpenStudy (anonymous):

I don't understand the problem?

OpenStudy (anonymous):

you have to find the area of the shaded region of the circle above

OpenStudy (anonymous):

sorry i forgot to shade it in when i drew it.

OpenStudy (anonymous):

Btw I'll become anyone who helps me's fan.

OpenStudy (anonymous):

Do you have any laws to use?

OpenStudy (blockcolder):

The area of the shaded region is \[\pi r^2\]Your r is 20, so r^2=400. Now, pi is approx. 3.1, so the approximate area is 400*3.1=1240 cm^2.

OpenStudy (anonymous):

Why pi = 3.1 ? blockcolder?

OpenStudy (anonymous):

I know that pi = 22/7 or 3.14

OpenStudy (anonymous):

yes, its 3.14

OpenStudy (blockcolder):

It's just an approximation I used to make the multiplication easier. :D

OpenStudy (anonymous):

I think you have to divide the total area by 4 because it's quarter of a circle not the whole circle

OpenStudy (anonymous):

I have a question, isnt pi r 2 the formula for the area of the circle?

OpenStudy (blockcolder):

Oh. I thought the shaded area was the whole circle. :P

OpenStudy (anonymous):

No

OpenStudy (anonymous):

no, just were it says 20 cm

OpenStudy (blockcolder):

Right. So you need to divide my answer by 4.

OpenStudy (anonymous):

pi r 2 / 4?

OpenStudy (anonymous):

But I'm not sure if this area is a quarter of the circle or what?

OpenStudy (anonymous):

yes, a quarter.

OpenStudy (anonymous):

1240/4 = 314.28 cm2

OpenStudy (blockcolder):

Yes, that one. :D

OpenStudy (anonymous):

it was like this.

OpenStudy (anonymous):

So yes that's right horsesrock

OpenStudy (anonymous):

ok, thanks, smr.

OpenStudy (anonymous):

do i write the whole thing, like you typed it?

OpenStudy (anonymous):

314.28?

OpenStudy (anonymous):

yes you have to write the whole area and then divide it by 4

OpenStudy (anonymous):

ok. thanks.

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

For helping me.

OpenStudy (anonymous):

'm going to"fan" you.

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

:D as a "thank you".

OpenStudy (anonymous):

no prob ;D

OpenStudy (anonymous):

I just dif, as a matter of fact.

OpenStudy (anonymous):

did*

OpenStudy (anonymous):

I will also write a testimonial

OpenStudy (anonymous):

ok great

OpenStudy (anonymous):

can you message me?